# Scaling down Deep Learning

### v2

@@ 4 Example use cases
Metalearning an activation function. Having implemented a “minimal working example” of gradient-based metalearning, we realized that it permitted a simple and novel extension: metalearning an activation function. With a few more hours of researcher time, we were able to parameterize our classifier’s activation function with a second neural network and then learn the weights using meta-gradients. As Figure 6 shows, our learned activation function substantially outperforms baseline nonlinearities such as ReLU, Elu, and Swish [^6] Run in browser: bit.ly/38V4GlQ (Clevert et al., 2016; Ramachandran et al., 2018).
@@ Acknowledgements
Thanks to Luke Metz for all the fun conversations. Thanks to Tony Zador for encouraging me to finally publish this.
Appendix A Hyperparameters

### v3

@@ 4 Example use cases
Metalearning an activation function. Having implemented a “minimal working example” of gradient-based metalearning, we realized that it permitted a simple and novel extension: metalearning an activation function. With a few more hours of researcher time, we were able to parameterize our classifier’s activation function with a second neural network and then learn the weights using meta-gradients. As Figure 6 shows, our learned activation function substantially outperforms baseline nonlinearities such as ReLU, Elu, and Swish. [^6] Run in browser: bit.ly/38V4GlQ We note that previous works (Clevert et al., 2016; Ramachandran et al., 2018; Vercellino and Wang, ) have tried to optimize activation functions, but none have done so with analytical gradients computed via bilevel optimization.
@@ Acknowledgements
Thanks to Luke Metz for all the fun conversations. Thanks to Tony Zador for encouraging me to finally publish this. Thanks to Dmitry Kobak for helping create the tSNE plots comparing MNIST and MNIST-1D shown in Figure 10.
Appendix A Supplementary figures
Dimensionality reduction. We used tSNE to reduce the dimensionality of MNIST and MNIST-1D so as to plot the datasets in two dimensions. We show the results in Figure 10 with each example colored according to its class label. We observe ten well-defined clusters in the MNIST dataset, suggesting that most examples in most classes are linearly separable from one another. By contrast, we observe few well-defined clusters in the MNIST-1D dataset, suggesting that a classifier must learn a nonlinear representation of the data in order to separate the classes properly. Thanks to Dmitry Kobak for helping with this visualization.
Figure 10: Visualizing the MNIST and MNIST-1D datasets with tSNE. The well-defined clusters in the MNIST plot indicate that the majority of the examples are linearly separable according to class. The MNIST-1D plot, meanwhile, reveals a lack of well-defined clusters which suggests that nonlinear features are much more important for successful classification.
Appendix B Hyperparameters